# Using a Deep Learning Model to Simulate Human Stock Traders’ Methods of Chart Analysis

### 30 Dec 2023 v2

@@ 4 Test Results
Table 9: Korea backtest result on test period
@@
Table 10: US backtest result on test period
Backtest results from year 2020 to 2022, averaged with 5 runs are shown in table 9 and 10 order by the total return. The market returns are 2.81%, 0.78%, for KOSPI, KOSDAQ (Korea market) and 17.85%, 15.12%, 60.99% for S&P500, NASDAQ, NYSE, AMEX (US market) during this period. The labels return varies widely according to periods. Only label_5_tp10_ls10_threshold_0.99_sidecut_False on Korea market and label_20_tp20_ls20_threshold_0.0_sidecut_True on Us market gives 40% excess return respect to best random market result on both validation and test period.
@@ 4.1 Example Charts Model Buys
Stock charts that model buy are shown in Fig 3 and 4. Stock bar chart of 600 days that model predicts a 10% rise in price and actually rises are represented. The Y-axis represents the stock price, measured in Korean Won for Korea stocks and in dollars for US stocks. The X-axis indicates the ordinal position of the input data. The top left corner displays the predicted date. Blue indicates a bullish candlestick and gray indicates a bearish. OHLCV candles are shown with 5-day SMA (Simple Moving Average), 20/60/120/240/480-day EMA (Exponential Moving Average).
The model buys stocks with various chart patterns, including sideways, uptrend, and downtrend stocks. 3 shows an example of a Korea stock that broke through the resistance area of 200-300 days with a long white body candle. It looks like the stock price has risen a lot, but the model expects it to rise additionally, which it does. 4 shows an example of US stock which the price went down in consecutive days while the price-rising trend is formed in the long term. The price bounces off the 480-day moving average and achieves 10% rise in price in next few days not shown in the chart. The accuracy of the model varies according to the market condition. Above patterns work well when the market is bullish, however fails rapidly if the market is bearish. One can further consider avoiding these high momentum stocks.
@@ 5 Discussions and limits
Slippage was not taken into account. Slippage in trading refers to the difference between the expected and actual execution prices of a trade, occurring due to market volatility, liquidity, or order size. All the stocks were assumed to be sold at exact 10% price. This can be a problem for trading low-volume stocks. Stock gaps were not taken into account, the same fixed profit or loss was taken even when the day’s open price was formed above or below the fixed percentage of 10%. In real-life trading, the investor may get profit/loss more than the specified percentage. Delisted tickers were not taken into account. About 2% to 3% of Korean and US stocks are delisted annually. This percentage might not be few, but this risk can be reduced by abstaining from trading during the applicable compliance period. The model was trained using the closing price adjusted after after-hour trades, which means that the model buys stocks at the exact close price at the end of the market. In a real trading scenario, one might have to buy before the market end or place an order the next day at the previous day’s close price.
6 Future Work
Naively labeling the stock charts by whether it reaches the given percentage in given period may contain many noises. The fact that label_1_tp10_ls10 doesn’t give the best total return although it gives the best F1 macro averaged score for both Korea and US market might be the evidence. We plan to initially select the charts that shows traditional trading patterns like pullback or breakthrough patterns and label them whether they end in profit or loss, and feed this information into the model to learn the specific conditions for the strategy’s success.
@@ 7 Conclusion
In this paper, a deep learning model is used to simulate how professional technical analysts trade by looking at the stock charts. Three labels were assigned according to whether a stock increased or decreased by equal to or more than 10% and whether the stock had less than 10% change in price within the given days. A Resnet model with a window size of 600 was used to predict the labels. A sample weighted loss function and thresholding, which allows the trading of stocks that has softmax logit value larger than a threshold, were deployed to increase the probability of the prediction. Backtesting on years 2020 to 2022 shows that the proposed approach in this paper generates a profit of more than 49% on the Korea market, and more than 40% on the US stock market. As far as we know, this paper is the first to utilize softmax logits of deep learning models in trading schemes with recent backtest results that far exceeds the market return. And this study suggests that setting a 5-day stop-loss period is advantageous for existing small-scale retail chart traders.”

### 08 Apr 2024 v3

@@ 4 Test Results
4.1 Comparison with Market Index
The model is backtested on year 2020 to 2022 for test period. The Return (%), maximum drawdown (%), and Sharpe ratio of US and Korea’s market index in shown on table 8. For Korea market, our best model (model_label_20_tp10_ls10_sidecut_False) far exceeds the total return by 72% and the Sharpe ratio by 1.42. However, for US market our best model (model_label_30_tp10_ls10_sidecut_False) fail to exceed total return and Sharpe ratio of the AMEX index although it gives 10% more total return and 0.37 more Sharpe ratio than investing in S&P500 and NASDAQ index. We think this is because the AMEX market performed extremely well during this test period. On validation period, AMEX ranks last compared to our model and other benchmarks in the validation period 9. Our results far exceeds the Sharpe ratio of all the market index although it may just be a result of over-fitting.
[…]
market_stats.csv
Table 8: Return of Korea and US Market Indices during the Test Period
[…]
val_market_stats.csv
Table 9: Return of Korea and US Market Indices during the Validation Period
4.2 Comparison of Quarterly Return of Market Index
We compare the quarterly return of our best model (model_label_20_tp10_ls10_sidecut_False for Korea market, and model_label_30_tp10_ls10_sidecut_False for US market) with the market return. The quarterly returns are calculated using one of the five random runs with the equity curve of the portfolio. The result table and bar plot of three month return during the test period is shown on table 10, 11 and figure 6, 5. Although our model does not consistently beat the market on quarterly basis, big losses are minimized when they occur.
Figure 5: Barplot of Quarterly return of KR market during the test period
Figure 6: Barplot of Quarterly return of US market during the test period
[…]
kr_quarterly.csv
Table 10: Quarterly return of KR market during the test period
[…]
us_quarterly.csv
Table 11: Quarterly return of US market during the test period
4.3 Comparison with the Randomly Chosen Stocks and Market Index
Our dataset includes survival bias, which means that stocks that are deslited are not included in historical prices. This can result in higher returns and Sharpe ratios, as surviving stocks are likely to have performed well. However preprocessed datasets that takes these into account are expensive. Hence, we compare the trained model with results of holding randomly selected stocks until they hit a predetermined percentage limit over the given period. These are shown with prefix ’random_’ in front of the label. Box plot of total return of five runs are shown on figure 7, 8. Top ten result when ordered by Sharpe ratio is shown on table 12, 13. If the model selects stocks that are most likely to rise, it should yield better returns than randomly selecting stocks excluding de-listed ones. Despite the stocks are randomly selected, some number of backtests outperform the market return due to high turnover rate, which has the effect similar to investing higher amount of initial capital because when a stock is sold another stock is purchased immediately with the available cash. Trained model generally tends to perform well than randomly chosen stocks, as indicated by blue label. For US market model_label_30_tp20_ls20_sidecut_False ranks the best consistent with the validation result. For Korean market model_label_20_tp10_ls10_sidecut_False ranks fourth when ordered by Sharpe ratio, however it gives the best total return and still far exceeds the randomly chosen market result. Although gives the best total return on both valdation and test period, when ordered by Sharpe ratio. However both results far exceeds the best random result.
Figure 7: total return of Korea market during the test period
@@ 4.3 Comparison with the Randomly Chosen Stocks and Market Index
Table 12: KR market test period backtest result top 10 ordered by Sharpe ratio
Figure 8: total return of US market during the test period
@@
Table 13: US market test period backtest result top 10 ordered by Sharpe ratio
4.4 Comparison with Other Work
We compare other work with another labeling scheme on stock price. Han et al. (2023) test the effect of N-period-min-max (NPMM) labeling with technical indiators and XGBoost algorithm. In their study, they assigned labels (0 or 1) based on whether the current price reached the lowest or highest price within a predefined window. While their model achieved profitability across 92 NASDAQ stocks, their evaluation focused on individual stock performance and did not address portfolio rebalancing strategies (what proportion of original cash to enter each trade). When labeled with NPMM method very few points have labels,resulting in insufficient data points for XGBoost training on certain stocks. To compare with our result, we train XGBoost model on 84 tickers specified in the paper (8 tickers were excluded because data was not available) by setting t he train period from 2009-01-02 to 2019-12-31. The model was backtested on 2020-01-02 to 2022-12-31. NPMM labeling period of 61 days was used, which was reported to have based performance in the paper. Entry ratio was tested with 0.05, 0.1, 0.2. The result which averages five runs are summarized on table 14, and figures 9, 10
[…]
npmm_bt_result.csv
Table 14: NPMM Backtest Results compared with ours
Figure 9: total return of NPMM compared with Ours
Figure 10: Sharpe ratio of NPMM compared with Ours
When trading with NPMM labeling, only about 19 trades occurred giving return of 4.79%, having Sharpe ratio of 0.19 and Maximum Drawdown of 17.35%. Our best model far exceeds this with return of 35.54%, having Sharpe ratio of 0.61 and Maximum Drawdown of 27.17%.
4.5 Effect of Thresholding on Backtest Result
The boxplots illustrating Sharpe ratios regarding different thresholds of the best model for KR market is shown on Figure 11 and and US market on Figure 12.
For test period, sharpe ratio tends to increase as you increase threshold, showing a J-like curve. Beyond a certain threshold, the model ceases to provide outputs due to none of the softmax logits surpassing the specified threshold. Because there were big downfall of stock market during the test period, it seems it was advantageous to only buy few stocks that were certain.
Figure 11: Box Plot of Sharpe Ratio with different thresholds during test period using KR Best Model
Figure 12: Box Plot of Sharpe Ratio with different thresholds during test period using US Best Model
@@ 4.6 Example Charts Model Buys
Stock bar chart of 600 days that model predicts a 10% rise in price and actually rises are represented in Fig 13 and 14. The Y-axis represents the stock price, measured in Korea Won for Korea stocks and in dollars for US stocks. The X-axis indicates the ordinal position of the input data. The top left corner displays the predicted date. Blue indicates a bullish candlestick and gray indicates a bearish. OHLCV candles are shown with 5-day SMA (Simple Moving Average), 20/60/120/240/480-day EMA (Exponential Moving Average).
The model buys stocks with various chart patterns, including sideways, uptrend, and downtrend stocks. 13 shows an example of a Korea stock that broke through the resistance area of 200-300 days with a long white body candle. It looks like the stock price has risen a lot, but the model expects it to rise additionally, which it does. 14 shows an example of US stock which the price went down in consecutive days while the price-rising trend is formed in the long term. The price bounces off the 480-day moving average and achieves 10% rise in price in next few days not shown in the chart. The accuracy of the model varies according to the market condition. Above patterns work well when the market is bullish, however fails rapidly if the market is bearish. One can further consider avoiding these high momentum stocks.
@@ 5 Discussions and limits
We notice that although label_1_tp10_ls10 gives the highest F1 macro score for both Korea and Us labels, it doesn’t given necessarily gives the best total return. This seems to be do to over-fitting, as there may be no underlying reasons that makes us possible to predict the stocks that will rise 10% in that day and the model just fits to noise signals. Slippage was not taken into account. Slippage in trading refers to the difference between the expected and actual execution prices of a trade, occurring due to market volatility, liquidity, or order size. All the stocks were assumed to be sold at exact 10% price. This can be a problem for trading low-volume stocks. Stock gaps were not taken into account, the same fixed profit or loss was taken even when the day’s open price was formed above or below the fixed percentage of 10%. In real-life trading, the investor may get profit/loss more than the specified percentage. Delisted tickers were not taken into account. About 2% to 3% of Korea and US stocks are delisted annually. This percentage might not be small, but this risk can be reduced by abstaining from trading during the applicable compliance period. The model was trained using the closing price adjusted after after-hour trades, which means that the model buys stocks at the exact close price at the end of the market. In a real trading scenario, one might have to buy before the market end or place an order the next day at the previous day’s close price.
@@ 6 Conclusion
In this paper, a deep learning model is used to simulate how professional technical analysts trade by looking at the stock charts. Three labels were assigned according to whether a stock increased or decreased by equal to or more than 10% or 20% and whether the stock had less than 10% or 20% change in price within the given days. A Resnet model with a window size of 600 was used to predict the labels. To increase the probability of the prediction, only stocks that has softmax logit value larger than a chosen threshold were traded. Backtesting on years 2020 to 2022 shows that the proposed approach in this paper outperforms index benchmarks. On Korea market it gave return of 75.36% having Sharpe ratio of 1.57, which far exceeds the market return by 36% and 0.61. On US market it gives total return of 35.54% with Sharpe ratio of 0.61, which outperforms other benchmarks such as NASDAQ, S&P500, DOW JONES index by 17.69% and 0.27, although it fails to outperform AMEX index. As far as we know, this paper is the first to utilize softmax logits of deep learning models in trading schemes with recent backtest results that far exceeds the market return.